OpenStudy (anonymous):

Please help!!! I would like to know what I am doing wrong, that my answer is different!! Let C(x) =9x +25 be the cost to produce x units of a product, and let R(x)=-x^2+31x be the revenue. Find the maximum profit. Why answer is 11? Thnaks!! :)

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

I got 12

OpenStudy (anonymous):

is this calculus or quadratic equations

OpenStudy (anonymous):

Is Calculus.

OpenStudy (anonymous):

maximize R(x) -C(x)

OpenStudy (anonymous):

so your derivative is...?

OpenStudy (anonymous):

-2x +........?

OpenStudy (anonymous):

-2x+50+25?? I think this is the part that is wrong

OpenStudy (anonymous):

hmm yeah... R-C = -x^2 +22x +25

OpenStudy (anonymous):

Ooh noo Sorry!!

OpenStudy (anonymous):

whoops, should be: R-C = -x^2 +22x -25 not that it matters much for our purposes...

OpenStudy (anonymous):

so the derivative is -2x + 22

OpenStudy (anonymous):

set it equal to zero and solve for x

OpenStudy (anonymous):

that's your max.

OpenStudy (anonymous):

I'm confused now, where do you get the 22 from?

OpenStudy (anonymous):

derivative of 22x is 22

OpenStudy (anonymous):

may be u should revise your notes on derivative properties before attemptig these word problems

OpenStudy (anonymous):

cleared up @rocal2 ? questions?

OpenStudy (anonymous):

I am still trying to figure out why 22x?

OpenStudy (anonymous):

-x^2 +31x -( 9x +25)

OpenStudy (anonymous):

Yes, I got that part.. then next step: You substract 31x -(9x+25)

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=maximize+-x^2%2B31x+-+9x+-25 Profit = Revenue - Cost

OpenStudy (experimentx):

profit is not 11, profit is 96 at x=11 http://www.wolframalpha.com/input/?i=-x^2%2B31x+-+9x+-25%2F.x-%3E11

OpenStudy (anonymous):

I think they're asking for the number of units in order to maximize profit...

OpenStudy (anonymous):

@experimentX I understand that, but I gonna have a test and I won't allow to use the internet, how do i get the 11? thats my question. ;))

OpenStudy (anonymous):

find R-C we did that... what was it?

OpenStudy (anonymous):

@Algebraic! Yes they want the Maximum profit.

OpenStudy (anonymous):

R-C= -x^2+22x+25

OpenStudy (anonymous):

no they don't. not if the answer is 11. You're not reading the question properly or you're answering the wrong part of the question. anyway... right now we're finding the UNITS to be made so that profits are maximized...

OpenStudy (anonymous):

R-C= -x^2+22x-25 now... find the location of the max by taking the derivative and setting it equal to zero

OpenStudy (experimentx):

well ... i too suppose so.

OpenStudy (anonymous):

-2x+22=0 ??

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

Ok, good!! I undertand now I have to solve. But, the 22 can you explain me what you did?

OpenStudy (anonymous):

" -x^2 " what's the derivative?

OpenStudy (anonymous):

" 22x " what's the derivative?

OpenStudy (anonymous):

So propably that is why i am lost!! I have to double check my notes.

OpenStudy (anonymous):

you should be pretty familiar with straightforward and simple derivatives by now... especially if you're already doing optimization. definitely going to want to review the basic rules of differentiation for the test.

OpenStudy (anonymous):

\[(x ^{n}) ' =n*x ^{n-1}\] \[(c*x) ' = c\]

OpenStudy (anonymous):

etc.

OpenStudy (anonymous):

Ok, Yes that's true is an online class, and I have to teach myself. I will definately to go back and check... Thanks!